OpenStudy (anonymous):

find 2 vectors orthogonal to (0,1,1) and to each other

OpenStudy (regan):

[1,1,0] and [1,0,1]

OpenStudy (anonymous):

but the dot product of (1,0,1) and (0,1,1) is 1, not 0

OpenStudy (anonymous):

Your questions don't same to have enough info to work with. Where are you getting these questions? from a book? making them up for practice?

OpenStudy (anonymous):

(0,-1,1) should work

OpenStudy (anonymous):

a book

OpenStudy (anonymous):

So share with us. How did you determine <0,-1,1>

OpenStudy (anonymous):

wouldn't (0,1,1)*(0,-1,1) = 0?

OpenStudy (anonymous):

It may very well be. I am asking how did you find it?

OpenStudy (anonymous):

i just did it in my head

OpenStudy (anonymous):

So if you were instructing someone, you would tell them make a guess and then test it to see if it so?

myininaya (myininaya):

you saay let vector v=(v1,v2,v3) such that u dot v=0*v1+1*v2+1*v3=0 any vector v that satisfies that equation is orthogonal to (0,1,1)